The manager of a pharmacy wants to know if prescriptions are filled uniformly over the 7 days of the week. The manager takes a simple random sample of 245 prescription receipts and finds that they are distributed as follows:

Day

Monday

Tuesday

Wednesday

Thursday

Friday

Saturday

Sunday

Prescriptions

42

31

33

29

45

44

21

Which of the following is the appropriate null hypothesis for this test?

  1. H0: p1= p2= p3= p4= p5= p6= p7= 1/7
  2. H0: p1= p2= p3= p4= p5= 5/7 and p6= p7= 2/7
  3. H0: p1= 0.17, p2= 0.13, p3= 0.13, p4= 0.12, p5= 0.18, p6= 0.18, p7= 0.09
  4. None of the above

Under the null hypothesis of a uniform distribution of prescriptions over the 7 days of the week, the expected count of prescriptions for Monday is _____________ (show calculation).

Under the null hypothesis of a uniform distribution of prescriptions over the 7 days of the week, the chi-square contribution for Monday is _________________ (show calculation).

Under the null hypothesis of a uniform distribution of prescriptions over the 7 days of the week, the degrees of freedom for the chi-square test is _______________.
Answer #1

Null hypothesis:

H0: p1= p2= p3= p4= p5= p6= p7= 1/7

Under the null hypothesis of a uniform distribution of prescriptions over the 7 days of the week, the expected count of prescriptions for Monday is =np=245*1/7 =35

Under the null hypothesis of a uniform distribution of prescriptions over the 7 days of the week, the chi-square contribution for Monday is =(O-E)^2/E =(42-35)^2/35 =1.4

Under the null hypothesis of a uniform distribution of prescriptions over the 7 days of the week, the degrees of freedom for the chi-square test is =categories -1 =7-1 =6